# On circumcentered direct methods for monotone variational inequality problems

### v1

@@ 3.1 The algorithm
3.2 Convergence Analysis
We need some preliminary propositions for our convergence theorem.
@@ 3.2 Convergence Analysis
Suppose that $C=\cap_{i=1}^{m}C_{i}$. Let $V\subset\mathds{R}^{n}$ be a nonempty compact and convex set. Define ${\mathbf{V}}=:V\times\cdots\times V\subset\mathds{R}^{nm}$. Then, there exist $\bar{\tau}\in[0,1)$ such that
${\rm dist}(T_{{\mathbf{S}}_{k}\cap{\mathbf{X}},{\mathbf{D}}}(x),C)\leq\bar{\tau}\,{\rm dist}(x,C),$ (3.6)
for all $x\in V\setminus C$, where $T_{{\mathbf{S}}_{k}\cap{\mathbf{X}},{\mathbf{D}}}$ is as (1.21) and ${\mathbf{S}_{k}}=S_{k}^{1}\times\dots\times S^{m}_{k}$, where $S^{i}_{k}=S_{i}(x^{k})$ and $S^{i}:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ is a separating set for $C_{i}$ in the sense of Definition 1.2.
@@
${\rm dist}(T_{\mathbf{S}_{k}\cap\mathbf{V},\mathbf{D}}(x),C)\leq{\rm dist}((P_{\textbf{D}}\circ P_{\mathbf{S}_{k}\cap\mathbf{V}})(x),C),$ (3.7)
according to Proposition 2.6 with $K=\mathbf{S}_{k}\cap\mathbf{V}$ and $U=\mathbf{D}$. Now we invoke Definition 1.2(iv) for establishing that there exists $\tau_{i}\in(0,1]$ such that
${\rm dist}(P_{S_{i}\cap V}(x),C)\leq\tau_{i}\,{\rm dist(x,C)}$ (3.8)
@@
$\begin{split}{\rm dist}((P_{\textbf{D}}\circ P_{\mathbf{S}_{k}\cap\mathbf{V}})(x),C)&={\rm dist}\left(\frac{1}{m}\sum_{i=1}^{m}P_{S_{i}\cap V}(x),C\right)\\ &\leq\frac{1}{m}\sum_{i=1}^{m}{\rm dist}(P_{S_{i}\cap V}(x),C)\\ &\leq\frac{1}{m}\sum_{i=1}^{m}\tau_{i}\,{\rm dist(x,C)}\\ &\leq\bar{\tau}\,{\rm dist}(x,C),\end{split}$ (3.9)
for all $x\in V\setminus C$. The first inequality in (3.9) holds by convexity the distance function, the second one holds by (3.8) and the third one by the definition of $\bar{\tau}$. Using (3.7) and (3.9), we complete the proof. ∎
@@
Let $C=\cap_{i=1}^{m}C_{i}$. Consider the sequence $\{x^{k}\}$ defined by Algorithm 1. If $x^{k+1}=x^{k}$ for some $k$, then $x^{k}\in{\rm SOL(F,C)}$.
@@
We suppose that $x^{k+1}=x^{k}$, so in the product space $\mathds{R}^{nm}$ we have $\mathbf{x}^{k+1}=\mathbf{x}^{k}$. Since $\mathbf{x}^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)$, we get from Proposition 2.3 that $\mathbf{x}^{k}\in\mathbf{H}^{k}\subset\mathbf{S}^{k}\subset\mathds{R}^{nm}$. Therefore, $x^{k}\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. Given $j\in\{1,\ldots,m\}$, we have two cases according to the definition of $S_{i}^{k}$ in (2.4). If $S_{i}^{k}=C_{i}$, then we have $x^{k+1}=x^{k}\in C_{j}$. Otherwise, we have $g_{j}(x^{k})\leq g_{j}(x^{k})+\langle v^{jk},x^{k+1}-x^{k}\rangle\leq 0$, i.e., $x^{k}\in C_{j}$. By the above argument above, we conclude that for any $i\in\{1,\ldots,m\}$, we have $x^{k}\in C_{i}$. Therefore, $x^{k}\in C=\cap_{i=1}^{m}C_{i}$. Moreover, since $x^{k+1}$ is given by (3.3), using Proposition 2.1 (i) in the product space, with $x=\mathbf{x}^{k}-\frac{\beta_{k}}{\eta_{k}}T_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{x}^{k})$, the approximate circumcenter operator coincides with the projection operator, in view of Proposition 2.3, i.e., $T_{\mathbf{S}_{k},\mathbf{D}}(x)=P_{\mathbf{S}_{k}\cap\mathbf{D}}(x)$, and $K=\mathbf{S}_{k}$. Hence we obtain (first in the product space)
$\left\langle\mathbf{x}^{k+1}-(\mathbf{x}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\mathbf{x}^{k})),\mathbf{z}-\mathbf{x}^{k+1}\right\rangle\geq 0~{}~{}\forall\mathbf{z}\in\mathbf{S}_{k},$ (3.10)
and then, in ${\mathbf{R}}^{n}$,
$m\left\langle x^{k+1}-(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})),z-x^{k+1}\right\rangle\geq 0~{}~{}\forall z\in S_{i}^{k},$ (3.11)
for all $i\in\{1,\ldots,m\}$. Taking $x^{k+1}=x^{k}$ in (3.11) and taking into account the facts that $\beta_{k}>0,\eta_{k}\geq 1$ for all $k$ and $z\in C=\cap_{i=1}^{m}C_{i}\subset S_{i}^{k}$ in Proposition 2.4, we get $\langle F(x^{k}),z-x^{k}\rangle\geq 0$ for all $z\in C$. Then, we conclude that $x^{k}\in{\rm SOL(F,C)}$. ∎
@@
Take $\hat{z}\in C:=\cap_{i=1}^{m}C_{i}$ and $W\subset{\mathbb{R}}^{n}$ as in Assumption 3.5. Let $\{x^{k}\}$ be a sequence generated by Algorithm 1. Choose $\lambda>0$ such that $\left\lVert x^{0}-\hat{z}\right\rVert\leq\lambda$ and $W\subset B(\hat{z},\lambda)$. Then
  • (i)
@@
Since $\hat{z}\in C$, we get from (3.42) that $\hat{z}=T_{\mathbf{S}_{k},\mathbf{D}}(\hat{z})$. Then, in view of (3.3) and Proposition 2.1(i), we obtain
$\begin{split}\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}&=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k}))-\hat{z}\right\rVert^{2}=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(\hat{z})\right\rVert^{2}\\ &\leq\left\lVert\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-\hat{z}\right\rVert^{2}\\ &=\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\frac{\beta_{k}^{2}}{\eta_{k}^{2}}\left\lVert F(x^{k})\right\rVert^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\hat{z}\rangle\\ &\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\hat{z}\rangle.\end{split}$ (3.12)
@@
  • (i)
    if $x^{k}\in W$, applying Cauchy-Schwartz inequality in (3.12), the definition of $\eta_{k}$ and the fact that $W\subset B(\hat{z},\lambda)$, we conclude that
    $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2}+2\frac{\beta_{k}}{\eta_{k}}\left\lVert F(x^{k})\right\rVert\left\lVert x^{k}-\hat{z}\right\rVert\leq\lambda^{2}+\beta_{k}^{2}+2\beta_{k}\lambda.$ (3.13)
  • (ii)
@@
    $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2},$ (3.14)
@@
  • (i)
    Take $\hat{z}$ and $W$ as in Assumption 3.5, $\lambda>0$ such that $\left\lVert x^{0}-\hat{z}\right\rVert\leq\lambda$ and $W\subset B(\hat{z},\lambda)$, and $\bar{\beta}>0$ such that $\beta_{k}\leq\bar{\beta}$ for all $k$ (such a $\bar{\beta}$ exists because $\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty$). Let $\sigma=\sum_{j=0}^{\infty}\beta_{j}^{2}$. Define $\bar{\lambda}:=(\lambda^{2}+\bar{\beta}\lambda+\sigma)^{\frac{1}{2}}$. We claim that
    $\{x^{k}\}\subset B(\hat{z},\bar{\lambda}).$ (3.15)
@@
    $\left\lVert x^{\ell(k)+1}-\hat{z}\right\rVert^{2}\leq\lambda^{2}+\beta_{\ell(k)}^{2}+2\beta_{\ell(k)}\lambda\leq\lambda^{2}+\bar{\beta}\lambda+\beta_{\ell(k)}^{2}.$ (3.16)
    Iterating the inequality in Lemma 3.8(ii), since $x^{j}\notin W$ for $j$ between $\ell(k)+1$ and $k-1$, we obtain
    $\left\lVert x^{k}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{\ell(k)+1}-\hat{z}\right\rVert^{2}+\sum_{j=\ell(k)+1}^{k-1}\beta_{k}^{2}..$ (3.17)
    Combining (3.16) and (3.17)
    $\begin{split}\left\lVert x^{k}-\hat{z}\right\rVert^{2}&\leq\lambda^{2}+2\bar{\beta}\lambda+\sum_{j=\ell(k)}^{k-1}\beta_{j}^{2}\\ &\leq\lambda^{2}+2\bar{\beta}\lambda+\sum_{j=0}^{\infty}\beta_{j}^{2}\\ &=\lambda^{2}+2\bar{\beta}\lambda+\sigma\\ &=\bar{\lambda}^{2}.\end{split}$ (3.18)
    Thus, $x^{k}\in B(\hat{z},\bar{\lambda})$, proving the claim, and hence $\{x^{k}\}$ is bounded. For $\{F(x^{k})\}$, use the boundedness of $\{x^{k}\}$ and Proposition 2.7(iii).
  • (ii)
@@
    $\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert\leq\frac{\beta_{k}}{\eta_{k}}\left\lVert F(x^{k})\right\rVert\leq\beta_{k},$ (3.19)
@@
    $T_{\mathbf{S}_{k}\cap\mathbf{X},\mathbf{D}}(x)=T_{\mathbf{S}_{k},\mathbf{D}}(x)$ (3.20)
@@
    ${\rm dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x),C)\leq\tilde{\mu}\,{\rm dist}(x,C),$ (3.21)
@@
    ${\rm dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}),C)\leq\tilde{\mu}\,{\rm dist}(x^{k},C),$ (3.22)
@@
    $\begin{split}{\rm dist}(x^{k+1},C)&\leq\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert+{\rm dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}),C)\\ &\leq\beta_{k}+\tilde{\mu}\,{\rm dist}(x^{k},C),\end{split}$ (3.23)
    using (3.23) and (3.21) in the second inequality. Therefore, using Proposition 2.9 with $v_{k}=\beta_{k}$, $\mu=\tilde{\mu}$ and $\varepsilon_{k}={\rm dist}(x^{k},C)$, we obtain
    $\lim_{k\to\infty}{\rm dist}(x^{k},C)=0,$ (3.24)
@@
    $\left\lVert x^{k+1}-x^{k}\right\rVert\leq\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert+\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})-x^{k}\right\rVert\leq\beta_{k}+{\rm dist}(x^{k},C).$ (3.25)
@@
If $F$ is paramonotone and SOL$(F,C)\neq\emptyset$, then all cluster points of any sequence $\{x^{k}\}$ generated by Algorithm 1 solve VIP$(F,C)$.
@@
$\gamma_{k}(x):=\langle F(x^{k}),x^{k}-x\rangle.$ (3.26)
Take any $\bar{x}\in{\rm SOL}(F,C)$. From Proposition 2.1(i) and the definition of $\eta_{k}$, we get that
$\begin{split}\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}&=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(\bar{x})\right\rVert^{2}\\ &\leq\left\lVert\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}\right)-\bar{x}\right\rVert^{2}\\ &=\left\lVert x^{k}-\bar{x}\right\rVert^{2}+\frac{\beta_{k}^{2}}{\eta_{k}^{2}}\left\lVert F(x^{k})\right\rVert^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\bar{x}\rangle\\ &\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\beta_{k}\left(2\frac{\gamma_{k}(\bar{x})}{\eta_{k}}-\beta_{k}\right).\end{split}$ (3.27)
First we prove that $\{x^{k}\}$ has cluster points in SOL$(F,C)$ Since $\{(x^{k}\},\{F(x^{k}))\}$ are bounded by Lemma 3.9, in view of Proposition 2.8 it suffices to prove that $\{\gamma_{k}(\bar{x})\}$ has a nonpositive cluster point. Assume that this is not true. Our assumption implies that $\{\gamma(\bar{x})\}$ must be bounded away from zero for large $k$, i.e., there exist $\bar{k}$ and $\rho>0$ such that $\gamma_{k}(\bar{x})\geq\rho$ for all $k\geq\bar{k}$. Since $\{F(x^{k})\}$ is bounded, there exists $\theta>1$ such that $\left\lVert F(x^{k})\right\rVert\leq\theta$ for all $k$. Therefore,
$\eta_{k}=\max\{1,\left\lVert u^{k}\right\rVert\}\leq\max\{1,\theta\}=\theta,$ (3.28)
@@
$\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\beta_{k}(2\rho-\beta_{k}),$ (3.29)
@@
$\bar{\rho}\beta_{k}\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}.$ (3.30)

### v2

@@ 3.1. The algorithm
We also remark that the use of stepsizes satisfying (1.22)-(1.23) (called sometimes vanishing or $\ell_{2}\setminus\ell_{1}$ stepsizes) is a consequence of two facts: we do not assume Lipschitz continuity of $F$, which precludes the option of an exogenous constant stepsize, and our weaker monotonicity assumptions on $F$ exclude the implementation of successful line-searches for the stepsize.
3.2. Convergence Analysis
For the convergence analysis of Algorithm 1, we assume in this subsection that $F$ satisfies Assumption 1.7. We start with some preliminary results.
@@ 3.2. Convergence Analysis
Suppose that $C=\cap_{i=1}^{m}C_{i}$. Let $V\subset\mathds{R}^{n}$ be a nonempty compact and convex set. Define ${\mathbf{V}}=:V\times\cdots\times V\subset\mathds{R}^{nm}$. Then, there exists $\bar{\tau}\in[0,1)$ such that
(3.2) $\operatorname*{dist}(T_{{\mathbf{S}}_{k}\cap{\mathbf{X}},{\mathbf{D}}}(x),C)\leq\bar{\tau}\,\operatorname*{dist}(x,C),$
for all $x\in V\setminus C$, where $T_{{\mathbf{S}}_{k}\cap{\mathbf{X}},{\mathbf{D}}}$ is as (1.18) and ${\mathbf{S}_{k}}=S_{k}^{1}\times\dots\times S^{m}_{k}$, where $S^{i}_{k}=S_{i}(x^{k})$ and $S^{i}:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ is a separating operator for $C_{i}$ in the sense of Definition 1.4.
@@
(3.3) $\operatorname*{dist}(T_{\mathbf{S}_{k}\cap\mathbf{V},\mathbf{D}}(x),C)\leq\operatorname*{dist}((P_{\textbf{D}}\circ P_{\mathbf{S}_{k}\cap\mathbf{V}})(x),C),$
according to Proposition 2.6 with $K=\mathbf{S}_{k}\cap\mathbf{V}$ and $U=\mathbf{D}$. Now we invoke Definition 1.4(iv) to establish that there exists $\tau_{i}\in(0,1]$ such that
(3.4) $\operatorname*{dist}(P_{S_{i}\cap V}(x),C)\leq\tau_{i}\,\operatorname*{dist}(x,C).$
@@
(3.5) $\begin{split}\operatorname*{dist}((P_{\textbf{D}}\circ P_{\mathbf{S}_{k}\cap\mathbf{V}})(x),C)&=\operatorname*{dist}\left(\frac{1}{m}\sum_{i=1}^{m}P_{S_{i}\cap V}(x),C\right)\\ &\leq\frac{1}{m}\sum_{i=1}^{m}\operatorname*{dist}(P_{S_{i}\cap V}(x),C)\\ &\leq\frac{1}{m}\sum_{i=1}^{m}\tau_{i}\,\operatorname*{dist}(x,C)\\ &\leq\bar{\tau}\,\operatorname*{dist}(x,C),\end{split}$
for all $x\in V\setminus C$. The first inequality in (3.5) holds by convexity of the distance function, the second one holds by (3.4) and the third one by the definition of $\bar{\tau}$. Using (3.3) and (3.5), we complete the proof. ∎
@@
Let $C=\cap_{i=1}^{m}C_{i}$. Consider the sequence $\{x^{k}\}$ defined by Algorithm 1. If $x^{k+1}=x^{k}$ for some $k$, then $x^{k}\in{\operatorname{SOL}(F,C)}$.
@@
We suppose that $x^{k+1}=x^{k}$, so in the product space $\mathds{R}^{nm}$ we have $\mathbf{x}^{k+1}=\mathbf{x}^{k}$. Since $\mathbf{x}^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)$, we get from Proposition 2.3 that $\mathbf{x}^{k}\in\mathbf{H}_{k}\subset\mathbf{S}_{k}\subset\mathds{R}^{nm}$. Therefore, $x^{k}\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. Given $i\in\{1,\ldots,m\}$, we have two cases according to the definition of $S_{i}^{k}$ in (2.4). If $S_{i}^{k}=C_{i}$, then we have $x^{k+1}=x^{k}\in C_{i}$. Otherwise, we have $g_{i}(x^{k})\leq g_{i}(x^{k})+\langle v^{k}_{i},x^{k+1}-x^{k}\rangle\leq 0$, where $v^{k}_{i}\in\partial g_{i}(x^{k})$ i.e., $x^{k}\in C_{i}$. By the above argument above, we conclude that for any $i\in\{1,\ldots,m\}$, we have $x^{k}\in C_{i}$. Therefore, $x^{k}\in C=\cap_{i=1}^{m}C_{i}$. Moreover, since $x^{k+1}$ is given by (3.1), using Proposition 2.1 (i) in the product space, with $x=\mathbf{x}^{k}-\frac{\beta_{k}}{\eta_{k}}T_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{x}^{k})$, the approximate circumcenter operator coincides with the projection operator, in view of Proposition 2.3, i.e., $T_{\mathbf{S}_{k},\mathbf{D}}(x)=P_{\mathbf{S}_{k}\cap\mathbf{D}}(x)$, and $K=\mathbf{S}_{k}$. Hence, we obtain (first in the product space)
(3.6) $\left\langle\mathbf{x}^{k+1}-(\mathbf{x}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\mathbf{x}^{k})),\mathbf{z}-\mathbf{x}^{k+1}\right\rangle\geq 0~~\forall\mathbf{z}\in\mathbf{S}_{k},$
and then, in $\mathds{R}^{n}$,
(3.7) $m\left\langle x^{k+1}-(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})),z-x^{k+1}\right\rangle\geq 0~~\forall z\in S_{i}^{k},$
for all $i\in\{1,\ldots,m\}$. Taking $x^{k+1}=x^{k}$ in (3.7) and taking into account the facts that $\beta_{k}>0,\eta_{k}\geq 1$ for all $k$ and $z\in C=\cap_{i=1}^{m}C_{i}\subset S_{i}^{k}$ in Proposition 2.4, we get $\langle F(x^{k}),z-x^{k}\rangle\geq 0$ for all $z\in C$. Then, we conclude that $x^{k}\in{\operatorname{SOL}(F,C)}$. ∎
@@
Take $\hat{z}\in C\coloneqq\cap_{i=1}^{m}C_{i}$ and $W\subset{\mathbb{R}}^{n}$ as in Assumption 1.9. Let $\{x^{k}\}$ be a sequence generated by Algorithm 1. Choose $\lambda>0$ such that $\left\lVert x^{0}-\hat{z}\right\rVert\leq\lambda$ and $W\subset B(\hat{z},\lambda)$. Then
  1. (i)
@@
[…]
@@
  1. (i)
    if $x^{k}\in W$, applying the Cauchy-Schwarz inequality in (3.8), the definition of $\eta_{k}$ and the fact that $W\subset B(\hat{z},\lambda)$, we conclude that
    (3.9) $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2}+2\frac{\beta_{k}}{\eta_{k}}\left\lVert F(x^{k})\right\rVert\left\lVert x^{k}-\hat{z}\right\rVert\leq\lambda^{2}+\beta_{k}^{2}+2\beta_{k}\lambda.$
  2. (ii)
@@
    (3.10) $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2},$
@@
  1. (i)
    Take $\hat{z}$ and $W$ as in Assumption 1.9, $\lambda>0$ such that $\left\lVert x^{0}-\hat{z}\right\rVert\leq\lambda$ and $W\subset B(\hat{z},\lambda)$, and $\bar{\beta}>0$ such that $\beta_{k}\leq\bar{\beta}$ for all $k$ (such a $\bar{\beta}$ exists because $\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty$). Let $\sigma=\sum_{j=0}^{\infty}\beta_{j}^{2}$. Define $\bar{\lambda}\coloneqq(\lambda^{2}+\bar{\beta}\lambda+\sigma)^{\frac{1}{2}}$. We claim that
    (3.11) $\{x^{k}\}\subset B(\hat{z},\bar{\lambda}).$
@@
    (3.12) $\left\lVert x^{\ell(k)+1}-\hat{z}\right\rVert^{2}\leq\lambda^{2}+\beta_{\ell(k)}^{2}+2\beta_{\ell(k)}\lambda\leq\lambda^{2}+\bar{\beta}\lambda+\beta_{\ell(k)}^{2}.$
    Iterating the inequality in Lemma 3.3(ii), since $x^{j}\notin W$ for $j$ between $\ell(k)+1$ and k1k-1, we obtain
    (3.13) $\left\lVert x^{k}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{\ell(k)+1}-\hat{z}\right\rVert^{2}+\sum_{j=\ell(k)+1}^{k-1}\beta_{k}^{2}.$
    Combining (3.12) and (3.13), we obtain
    (3.14) $\begin{split}\left\lVert x^{k}-\hat{z}\right\rVert^{2}&\leq\lambda^{2}+2\bar{\beta}\lambda+\sum_{j=\ell(k)}^{k-1}\beta_{j}^{2}\\ &\leq\lambda^{2}+2\bar{\beta}\lambda+\sum_{j=0}^{\infty}\beta_{j}^{2}\\ &=\lambda^{2}+2\bar{\beta}\lambda+\sigma\\ &=\bar{\lambda}^{2}.\end{split}$
    Thus, $x^{k}\in B(\hat{z},\bar{\lambda})$, proving the claim, and hence $\{x^{k}\}$ is bounded. For $\{F(x^{k})\}$, we use the boundedness of $\{x^{k}\}$, which implies boundedness of $\{F(x^{k})\}$, since $F$ is continuous on $\mathds{R}^{n}$ by Assumption (1.7).
  2. (ii)
@@
    (3.15) $\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert\leq\frac{\beta_{k}}{\eta_{k}}\left\lVert F(x^{k})\right\rVert\leq\beta_{k},$
@@
    (3.16) $T_{\mathbf{S}_{k}\cap\mathbf{X},\mathbf{D}}(x)=T_{\mathbf{S}_{k},\mathbf{D}}(x),$
@@
    (3.17) $\operatorname*{dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x),C)\leq\tilde{\mu}\,\operatorname*{dist}(x,C),$
@@
    (3.18) $\operatorname*{dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}),C)\leq\tilde{\mu}\,\operatorname*{dist}(x^{k},C),$
@@
    (3.19) $\begin{split}\operatorname*{dist}(x^{k+1},C)&\leq\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert+\operatorname*{dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}),C)\\ &\leq\beta_{k}+\tilde{\mu}\,\operatorname*{dist}(x^{k},C),\end{split}$
    using (3.19) and (3.17) in the second inequality. Therefore, using Proposition 2.9 with $v_{k}=\beta_{k}$, $\mu=\tilde{\mu}$ and $\varepsilon_{k}=\operatorname*{dist}(x^{k},C)$, we obtain
    (3.20) $\lim_{k\to\infty}\operatorname*{dist}(x^{k},C)=0,$
@@
    (3.21) $\left\lVert x^{k+1}-x^{k}\right\rVert\leq\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert+\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})-x^{k}\right\rVert\leq\beta_{k}+\operatorname*{dist}(x^{k},C).$
@@
If $F$ is paramonotone and $\operatorname{SOL}(F,C)\neq\emptyset$, then all cluster points of any sequence $\{x^{k}\}$ generated by Algorithm 1 solve $\operatorname{VIP}(F,C)$.
@@
(3.22) $\gamma_{k}(x)\coloneqq\langle F(x^{k}),x^{k}-x\rangle.$
Take any $\bar{x}\in\operatorname{SOL}(F,C)$. From Proposition 2.1(i) and the definition of $\eta_{k}$, we get that
(3.23) $\begin{split}\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}&=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(\bar{x})\right\rVert^{2}\\ &\leq\left\lVert\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}\right)-\bar{x}\right\rVert^{2}\\ &=\left\lVert x^{k}-\bar{x}\right\rVert^{2}+\frac{\beta_{k}^{2}}{\eta_{k}^{2}}\left\lVert F(x^{k})\right\rVert^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\bar{x}\rangle\\ &\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\beta_{k}\left(2\frac{\gamma_{k}(\bar{x})}{\eta_{k}}-\beta_{k}\right).\end{split}$
First we prove that $\{x^{k}\}$ has cluster points in $\operatorname{SOL}(F,C)$ Since $\{(x^{k}\},\{F(x^{k}))\}$ are bounded by Lemma 3.4, in view of Proposition 2.8 it suffices to prove that $\{\gamma_{k}(\bar{x})\}$ has a nonpositive cluster point. Assume that this is not true. Then, our assumption implies that $\{\gamma(\bar{x})\}$ must be bounded away from zero for large $k$, i.e., there exist $\bar{k}$ and $\rho>0$ such that $\gamma_{k}(\bar{x})\geq\rho$ for all $k\geq\bar{k}$. Since $\{F(x^{k})\}$ is bounded, there exists $\theta>1$ such that $\left\lVert F(x^{k})\right\rVert\leq\theta$ for all $k$. Therefore,
(3.24) $\eta_{k}=\max\{1,\left\lVert u^{k}\right\rVert\}\leq\max\{1,\theta\}=\theta,$
@@
(3.25) $\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\beta_{k}(2\rho-\beta_{k}).$
@@
(3.26) $\bar{\rho}\beta_{k}\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}.$